Evaluate the ∫sin³ θ cos θdθ


1
Expert's answer
2021-02-26T05:01:41-0500

"\int\sin^3\theta\cos\theta d\theta=\int\sin^3\theta d(\sin\theta)=\frac{1}{4}\sin^4\theta+C"